Non-Homogenous Linear Differential Equations

1. (2D ^ 2 + 3D + 1) * y = e ^ (- 3x)

2.(D^ 2 -2D+5)y=25x^ 2 +12

4. (D ^ 2 - 3D + 2) * y = 14sin 2x - 18cos 2x


1
Expert's answer
2022-02-14T17:03:04-0500

2D2 + 3D + 1)"\times" y =e(- 3x)


(D+1)(2D+1)=e(- 3x)


(D+1)(2D+1)="\frac{e}{3x}"


D=3, "\frac{1+\sqrt(2-3)}{3}"


y(x)= "\frac{1}{D^-2D^3+D-3}" ex


y(x)= "\frac{1}{1-2+1-2}" ex


y= "\frac{1}{1-2+1-2}" xex




(D2 -2D+5)y=25x2 +12

(D-1)(D-5)y=25x2+12


y= "\frac{25x2+12}{(D-1)(D-5)}"


D=1, "\frac{1+\sqrt(25+12)}{5}"


y(x)="\frac{6.08}{D^-2D^5+D-3}"




(D2 - 3D + 2)"\times" y = 14sin2x-18cos 2x


(D-1)(D+2)y=14sin2x-18cos2x


D=1, "\frac{sin2+\sqrt(14-2)}{5}"


y="\frac{(14sin^2(x))}{(d-1)(d+2)}" -"\frac{(18cos^2(x)}{(d-1)(d+2)}"